Suppose a company has $400 of fixed costs and a constant marginal cost of 10 cents....

  1. Suppose a company has $400 of fixed costs and a constant marginal cost of 10 cents. What are average total costs (ATC)

(a) Output of 10 units?

(b) Output of 100 units?

(c) Output of 1,000 units?
Answer #1

Answer

total cost function is

TC=FC+VC

FC=400 and VC=0.1Q

TC=400+0.1Q

ATC=TC/Q=400/Q+0.1

======

(a) Output of 10 units?

ATC(1)=400/10+0.1=40.1

(b) Output of 100 units?

ATC(100)=400/100+0.1=4.1

(c) Output of 1,000 units?

ATC(1000)=400/1000+0.1=0.5